A sample of 44 observations is selected from one population with a population standard deviation of 3.1. The sample mean is 101.0. A sample of 56 observations is selected from a second population with a population standard deviation of 5.0. The sample mean is 99.5. Conduct the following test of hypothesis using the 0.10 significance level.

H0 : μ1 = μ2

H1 : μ1μ2

  1. Is this a one-tailed or a two-tailed test?

  • One-tailed test

  • Two-tailed test

  1. State the decision rule. (Negative values should be indicated by a minus sign. Round your answers to 2 decimal places.)

  1. Compute the value of the test statistic. (Round your answer to 2 decimal places.)

  1. What is your decision regarding H0?

  • Reject H0

  • Do not reject H0

  1. What is the p-value? (Round your answer to 4 decimal places.)
Answer #1

n,=44 x= 101.0 0,=31 M2=56 X2=99.5 = 50 POPULATION STANDARD DEVIATION Xi = SAMPLE MEAN OF SAMI n - SAMPLE SIZES NULL HYPOTHESknown as two - sample 2- statistic 2=(5,-26? -(4,-M2) No.2 + 02² ni nz NORMAL which has standard 1. distribution (N,(0,1)) CA(101.o 2- Eteo -99.5) - (0) (3.12 (5.02 44 s6 2 = 15 = los d 0.2184 +0.446 No. 66482 6.81536 2= 1.8396 = 1.83 (9) WE HAVE TO> 20.os 2o-os ..21183 2-4/2 1.645 -1.645 REJECTION REGION FORHO ACCEPTANCE REGION FOR Ho REJECTION REGION FOR Ho ce pevalue T
Answer #2

Given that n = 44 J = 31 Ž, - 101.0 ₂ = 56 102=5.0 1x, = 99.5 | 2:0.10 Answer of Two tailed test (6) At 2 = 0.10 Lois (tworla